OpenStudy (anonymous):

use differences to find the pattern in the sequence. 3,9,13,19,31,53,89

OpenStudy (anonymous):

shoot i only have 5 minutes left to answer this question its my last

OpenStudy (anonymous):

Are you sure thats the right sequence?

OpenStudy (anonymous):

yeah it doesnt make sense right?

OpenStudy (anonymous):

??

OpenStudy (anonymous):

Yeah its kinda weird

OpenStudy (anonymous):

3*3=9+4=13+6=19........

OpenStudy (anonymous):

assuming that the pattern continues what should the eigth term be

OpenStudy (anonymous):

dont know sorry

OpenStudy (anonymous):

if u get the answer do post it,couldn't solve it..